why is change in enthalpy equal to heat absorbed by the system at constant pressure?

Grade:12th pass

1 Answers

Arun
5 years ago

If you see from the first law of thermodynamics,

del Q=dU+del W

where

Q is heat interaction, U is internal energy and W is the work done.

For a closed system,

del Q= dU+ PdV …..(1)

As we know from the definition of enthalpy,

H=U+PV

For change in enthalpy for constant pressure,

dH=dU+PdV ….(2)

So comparing the equations 1 and 2 we get,

del Q= dH

So, change in enthalpy is equal to heat interaction for a constant pressure process.

 

Regards

Arun (askIITians forum expert)